OpenStudy (anonymous):

Suppose the equation, y = 1600(1.03)x , represents the average cost of an education in a private college in 1992, where x stands for the number of years after 1992. The values in the table represent the first 4 years after 1992, rounded to the nearest dollar. Which is incorrect? Years after 1992 (x) 1 2 3 4 Cost of a private college education $ (y) 1,648 1,665 1,748 1,801 Answer 1 2 3 4

OpenStudy (anonymous):

@Shadowys

OpenStudy (anonymous):

A. 1 1648 B. 2 1665 C. 3 1748 D. 4 1801

OpenStudy (anonymous):

Whay do I simplify first?

OpenStudy (anonymous):

hmm...you just plug in x=1/2/3/4 into y=1600(1.03)^x and see if it's rite

OpenStudy (anonymous):

I got for my answer B. 2 1665

OpenStudy (anonymous):

@Shadowys

OpenStudy (anonymous):

Is it man? No wait A

OpenStudy (anonymous):

Yup A is my final answer correct bro?

OpenStudy (anonymous):

@Shadowys can you confirm?

OpenStudy (anonymous):

you sure? for x=1, y=1648.0 it's rite.

OpenStudy (anonymous):

So A is 1648 Yup Thanks (:

OpenStudy (anonymous):

:) Well the question asks for the incorrect number though.....